OpenStudy (anonymous):

Help With inverse functions 10 questions have answered 4 Please help medal to helper.

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

3

OpenStudy (anonymous):

f(2)=5 and g(5)=3

OpenStudy (anonymous):

@divu.mkr can you help me with others

OpenStudy (anonymous):

b(10)= 15 and m(15)= 46 so the answer is (C)

OpenStudy (anonymous):

its 2 f(4x+1/3) =3

OpenStudy (anonymous):

btw #3 is not 3 since its not one of the answers

OpenStudy (anonymous):

is it 2 for 3

OpenStudy (anonymous):

yes you are correct in q5

OpenStudy (anonymous):

5. correct

OpenStudy (anonymous):

okay is #3 2?

OpenStudy (anonymous):

ya the answer is 2 in 3

OpenStudy (anonymous):

its (c) 2-1/{0-(-2)}

OpenStudy (anonymous):

1

OpenStudy (anonymous):

divu whats your answer?

OpenStudy (anonymous):

its 1

OpenStudy (anonymous):

there is no 1 on this question what do you mean by 1

OpenStudy (anonymous):

actually i accidentally closed that tab in which i was typing... lol and i meant by option a

OpenStudy (anonymous):

@divu.mkr

OpenStudy (anonymous):

its y=7 coz its passing thru 7 ordinate

OpenStudy (anonymous):

sorry its x=7

OpenStudy (anonymous):

correction it is x=-4

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

so whats the answer lol

OpenStudy (anonymous):



OpenStudy (anonymous):

q9 its a

OpenStudy (anonymous):

so can eithe rof you help me with 5 worded problems? @divu.mkr @surjithayer

OpenStudy (anonymous):

for q.8. it is x=-4

OpenStudy (anonymous):

yeh we can y-y1=m(x-x1) m=2 y1=-3 and x1=1

OpenStudy (anonymous):

this is the first worded answer Please ill really appreciate it Thanks guys

OpenStudy (anonymous):

answer 2nd and 3rd both are equal to 16 3x+3=51 =>x=16 2(2x+1)=3(x+6) => x=16

OpenStudy (anonymous):

@divu.mkr Thanks for the help I really Appreciate it

OpenStudy (anonymous):

@surjithayer do you this question do you @divu.mkr

OpenStudy (anonymous):

\[30=\frac{ 4x-3 }{2 },60=4x-3, 4x=60+3=63,x=\frac{ 63 }{ 4 }\approx 16 weeks\]